# Simple homotopy types of even dimensional manifolds

### v2

@@ 5.2. Bass-Heller-Swan decomposition of $\operatorname{Wh}(C_{\infty}\times G)$
Let $G$ be a group. The decomposition of Theorem 5.9 restricts to the following isomorphisms:
@@
5.3. The groups $\mathcal{J}_{n}(C_{\infty}\times C_{m})$, $\mathcal{I}_{n}(C_{\infty}\times C_{m})$, and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$
@@
First note that by Proposition 5.10 and Theorem 5.7, for any group $G$ and any $n$, if $NK_{1}(\mathbb{Z}G)\neq 0$, then $\mathcal{J}_{n}(C_{\infty}\times G)$ and $\mathcal{I}_{n}(C_{\infty}\times G)$ are not finitely generated. If $G$ is finite and $n$ is even, then we also have the following.
@@
The next three theorems are the main results of Section 5. They will be established as a consequence of Theorems 11.1, 11.2 and 11.3, which will be proven in Section 11 (which is in Part III). These results were stated in the introduction as Theorems 1.6, 1.7 and 1.8 respectively.
@@
Suppose that $n$ is even and $m\geq 2$ is an integer.
@@
    $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=\infty$ if and only if $m$ is not square-free;
@@
Theorem 5.15.
Suppose that $n$ is even and $m\geq 2$ is an integer.
  1. (i)
  2. (ii)
    $|\mathcal{I}_{n}(C_{\infty}\times C_{m})|=\infty$ if and only if $m$ is not square-free;
  3. (iii)
    $|\mathcal{I}_{n}(C_{\infty}\times C_{m})|\to\infty$ super-exponentially in $m$.
Remark 5.16.
To write Theorem 5.14 (i) another way, the $m\geq 2$ for which $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=1$ are as follows:
$m=\begin{cases}\text{$p$},&\text{where $p\leq 19$ is prime}\\ \text{$2p$},&\text{where $p\leq 7$ is an odd prime}.\end{cases}$
Furthermore, Theorem 5.15 (i) states that $|\mathcal{I}_{n}(C_{\infty}\times C_{m})|=1$ if and only if $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=1$ or $m\in\{15,29\}$.
Proof of Theorems 5.14 and 5.15.
First note that Theorem 5.14 (ii) and Theorem 5.15 (ii) immediately follow from Theorem 5.8 and Lemma 5.12. For the proofs of (i) and (iii), we can therefore restrict to the case where $m$ is square-free, when $NK_{1}(\mathbb{Z}C_{m})=0$. By Propositions 5.10 and 5.13 we then have:
@@
The results follow from Theorems 11.1 and 11.2 (see Section 11). ∎
@@
$\liminf_{m\to\infty}|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|=1,\limsup_{m\to\infty}|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|=\infty.$
@@
Hence parts (i) and (iii) follows from Theorem 11.3, which will be proved in Section 11. Part (ii) follows from the fact that $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ is finite [Sw60] (see also Proposition 8.4 (ii)). ∎
@@ 6. Homotopy automorphisms of $S^{1}\times L$
Fix $k,m\geq 2$ and $q_{1},\ldots,q_{k}$ such that $\gcd(m,q_{j})=1$. Let $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ be a $(2k-1)$-dimensional lens space with $\pi_{1}(L)\cong C_{m}$. We will be interested in the product $S^{1}\times L$ and its homotopy automorphisms. Fix basepoints of $S^{1}$ and $L$ and let $i_{1}\colon S^{1}\rightarrow S^{1}\times L$ and $i_{2}\colon L\rightarrow S^{1}\times L$ denote the standard embeddings, we will identify $S^{1}$ and $L$ with their images in $S^{1}\times L$. Let $G=C_{\infty}\times C_{m}\cong\pi_{1}(S^{1}\times L)$.
6.1. The automorphism induced on the fundamental group
Our first goal is to determine which automorphisms of $G$ can be realised by homotopy automorphisms of $S^{1}\times L$, i.e. to describe the image of the map $\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(G)$. Note that the automorphisms of $G=C_{\infty}\times C_{m}$ can be expressed as matrices of the form $\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)$, where $a\in\{\pm 1\}$, $b\in C_{m}$ and $c\in C_{m}^{\times}\cong\operatorname{Aut}(C_{m})$.
Lemma 6.1 ([Co73, Statement 29.5]).
Let $c\in C_{m}^{\times}$. There is a homotopy automorphism $f\colon L\rightarrow L$ such that $\pi_{1}(f)=c$ if and only if $c^{k}\equiv\pm 1\,\,\text{mod}\,m$.
Lemma 6.2.
  1. (a)
    There is a diffeomorphism $f\in\operatorname{Diff}(S^{1}\times L)$ such that $\pi_{1}(f)=\left(\begin{smallmatrix}-1&0\\ 0&1\end{smallmatrix}\right)$.
  2. (b)
    There is a diffeomorphism $f\in\operatorname{Diff}(S^{1}\times L)$ such that $\pi_{1}(f)=\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)$.
Proof.
(a) Take $f=r\times\operatorname{Id}_{L}$ for an orientation-reversing diffeomorphism $r\colon S^{1}\rightarrow S^{1}$.
(b) Let $S^{2k-1}=\{(z_{1},\ldots,z_{k})\mid\sum|z_{j}|^{2}=1\}\subseteq\mathbb{C}^{k}$. Recall that $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ is the quotient of $S^{2k-1}$ by the $C_{m}$-action generated by $(z_{1},\ldots,z_{k})\mapsto(\zeta^{q_{1}}z_{1},\ldots,\zeta^{q_{k}}z_{k})$, where $\zeta=e^{2\pi i/m}$. Define $\phi\colon S^{1}\to\operatorname{Diff}(L)$ by $e^{2\pi it}\mapsto([z_{1},\ldots,z_{k}]\mapsto[\zeta^{q_{1}t}z_{1},\ldots,\zeta^{q_{k}t}z_{k}])$. Then we can define a suitable diffeomorphism $f$ by $f(x,y)=(x,\phi(x)(y))$. ∎
Lemma 6.3.
For every homotopy automorphism $f\colon S^{1}\times L\rightarrow S^{1}\times L$ there is a homotopy automorphism $\bar{f}_{2}\colon L\rightarrow L$ such that $f\circ i_{2}\simeq i_{2}\circ\bar{f}_{2}\colon L\rightarrow S^{1}\times L$.
Proof.
Since $\pi_{1}(L)\cong C_{m}$ is the torsion subgroup of $\pi_{1}(S^{1}\times L)\cong C_{\infty}\times C_{m}$, the restriction of $\pi_{1}(f)$ is an isomorphism $\pi_{1}(L)\rightarrow\pi_{1}(L)$, and hence $\operatorname{Im}\pi_{1}(f\circ i_{2})=\operatorname{Im}\pi_{1}(\rho)$ where the covering $\rho\colon\mathbb{R}\times L\rightarrow S^{1}\times L$ is the product of the universal covering of $S^{1}$ and $\operatorname{Id}_{L}$. By the lifting criterion [Ha02, Proposition 1.33], $f\circ i_{2}$ has a lift $\widetilde{f}_{2}\colon L\rightarrow\mathbb{R}\times L$ such that $f\circ i_{2}=\rho\circ\widetilde{f}_{2}$. Let $\pi\colon\mathbb{R}\times L\rightarrow L$ denote the projection. Since $\mathbb{R}$ is contractible, the universal covering map $\mathbb{R}\to S^{1}$ is null-homotopic. It follows that $\rho\simeq i_{2}\circ\pi$. Thus
$f\circ i_{2}=\rho\circ\widetilde{f}_{2}\simeq i_{2}\circ\pi\circ\widetilde{f}_{2}=i_{2}\circ\bar{f}_{2},$
where by definition $\bar{f}_{2}=\pi\circ\widetilde{f}_{2}\colon L\rightarrow L$.
This implies that $\pi_{1}(\bar{f}_{2})$ is the restriction of $\pi_{1}(f)$, so it is an isomorphism. Since $\pi_{j}(i_{2})$ is an isomorphism for $j>1$ and $f$ is a homotopy automorphism (hence $\pi_{j}(f)$ is an isomorphism), $\pi_{j}(\bar{f}_{2})$ is also an isomorphism. Therefore $\bar{f}_{2}$ is a homotopy automorphism too. ∎
Definition 6.4.
Let $A_{2k}(m)\leq\operatorname{Aut}(G)$ denote the subgroup of matrices $\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)$ such that $c^{k}\equiv\pm 1$ mod $m$.
Theorem 6.5.
We have $\operatorname{Im}(\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(G))=A_{2k}(m)$.
Proof.
Suppose that $f\colon S^{1}\times L\rightarrow S^{1}\times L$ is a homotopy automorphism with $\pi_{1}(f)=\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)$. By Lemma 6.3 there is a homotopy automorphism $\bar{f}_{2}\colon L\rightarrow L$ such that $f\circ i_{2}\simeq i_{2}\circ\bar{f}_{2}$. This implies that $\pi_{1}(\bar{f}_{2})=c$, so by Lemma 6.1 $c^{k}\equiv\pm 1$ mod $m$. Therefore $\pi_{1}(f)\in A_{2k}(m)$.
Let $c\in C_{m}^{\times}$ be such that $c^{k}\equiv\pm 1$ mod $m$. By Lemma 6.1 there is a homotopy automorphism $f\colon L\rightarrow L$ such that $\pi_{1}(f)=c$. Then $\operatorname{Id}_{S^{1}}\times f$ is a homotopy automorphism of $S^{1}\times L$ such that $\pi_{1}(\operatorname{Id}_{S^{1}}\times f)=\left(\begin{smallmatrix}1&0\\ 0&c\end{smallmatrix}\right)$. Matrices of this form, together with the matrices realised in Lemma 6.2, generate $A_{2k}(m)$. Therefore $A_{2k}(m)\subseteq\operatorname{Im}(\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(G))$. ∎
6.2. The Whitehead torsion
Next we prove that every homotopy automorphism of $S^{1}\times L$ is simple.
@@ 6.2. The Whitehead torsion
Remark 6.7.
When $m$ is odd, this also follows from results of Khan and Hsiang-Jahren. By [Kh17, Corollary 6.2], $\operatorname{hAut}(S^{1}\times L)$ is generated by maps of the form $f\times\operatorname{Id}_{L}$, $\operatorname{Id}_{S^{1}}\times g$ and $r_{h}$, where $f\in\operatorname{hAut}(S^{1})$, $g\in\operatorname{hAut}(L)$, $h\colon S^{1}\rightarrow\operatorname{Map}(L)$ and $r_{h}(x,y)=(x,h(x)(y))$. By Proposition 2.32, generators of the first two types have vanishing Whitehead torsion. If $m$ is odd, then it follows from [HJ83, Proposition 3.1] that $\tau(r_{h})=0$ (see also [Kh17, Corollary 6.4]).
The proof consists of two parts, which we will prove in the following two lemmas. For a CW complex $X$ and an integer $n\geq 0$, let $\operatorname{sk}_{n}(X)$ denote the $n$-skeleton of $X$. We will endow $S^{1}\times L$ with a standard CW decomposition, that we recall during the next proof.
@@
If $g\colon S^{1}\times L\rightarrow S^{1}\times L$ is a homotopy automorphism such that $g\big{|}_{L\cup\operatorname{sk}_{2k-2}(S^{1}\times L)}=\operatorname{Id}$, then $\tau(g)=0$.
@@
We will denote the generators of the components of $G$ by $\alpha\in C_{\infty}$ and $\beta\in C_{m}$. Let $\Sigma:=\sum_{j=0}^{m-1}\beta^{j}\in\mathbb{Z}C_{m}\subseteq\mathbb{Z}G$ be the norm element of $\mathbb{Z}C_{m}$.
The lens space $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ has a CW decomposition with one cell $E_{j}$ in each dimension $10\leq j\leq 2k-$ 1 (see e.g. [Co73, §28]). The corresponding cellular chain complex $(C^{L}_{j},d^{L}_{j})$ with $\mathbb{Z}C_{m}$ coefficients has $C^{L}_{j}\cong\mathbb{Z}C_{m}$ for every $j$ and
$d^{L}_{j}=\begin{cases}\Sigma&\text{if $j$ is even}\\ \beta^{r_{h}}-1&\text{if $j=2h-1$}\end{cases}$
where $r_{h}q_{h}\equiv 1$ mod $m$.
The product $M=S^{1}\times L$ has a CW decomposition containing the cells of $L$, plus a $(j+1)$-cell $E^{\prime}_{j+1}$, the product of $E_{j}$ and the $1$-cell of $S^{1}$, for every $10\leq j\leq 2k-$ 1. Its chain complex $(C^{M}_{j},d^{M}_{j})$ with $\mathbb{Z}G$ coefficients is the tensor product of $(C^{L}_{j},d^{L}_{j})$ and the chain complex of $S^{1}$ with $\mathbb{Z}C_{\infty}$ coefficients, $\mathbb{Z}C_{\infty}\stackrel{{\scriptstyle\alpha-1}}{{\longrightarrow}}\mathbb{Z}C_{\infty}$. It has $C^{M}_{0}\cong C^{M}_{2k}\cong\mathbb{Z}G$ and $C^{M}_{j}\cong\mathbb{Z}G^{2}$ for every $11\leq j\leq 2k-$ 1 (with $\mathbb{Z}G\oplus\left\{0\right\}$ corresponding to $E_{j}$ and $\left\{0\right\}\oplus\mathbb{Z}G$ corresponding to $E^{\prime}_{j}$). The differentials are
$\begin{gathered}d^{M}_{1}=\begin{pmatrix}\beta^{r_{1}}-1\\ \alpha-1\end{pmatrix},\quad d^{M}_{2h-1}=\begin{pmatrix}\beta^{r_{h}}-1&0\\ \alpha-1&\Sigma\end{pmatrix}\quad\text{for $h>1$},\\ d^{M}_{2k}=(1-\alpha,\beta^{r_{k}}-1),\quad d^{M}_{2h}=\begin{pmatrix}\Sigma&0\\ 1-\alpha&\beta^{r_{h}}-1\end{pmatrix}\quad\text{for $h<k$}.\end{gathered}$
We write elements of $\mathbb{Z}G^{j}$ as row vectors, so a homomorphism $\mathbb{Z}G^{j}\rightarrow\mathbb{Z}G^{h}$ is multiplication on the right by a $j\times h$ matrix.
We can assume that $g$ is cellular, let $g_{j}\colon C^{M}_{j}\rightarrow C^{M}_{j}$ denote the induced chain map. Since $g\big{|}_{L\cup\operatorname{sk}_{2k-2}(L\times S^{1})}=\operatorname{Id}$, we have $g_{j}=\operatorname{Id}$ for every $j\leq 2k-2$ and $g_{2k-1}\big{|}_{\mathbb{Z}G\oplus\left\{0\right\}}=\operatorname{Id}$.
We investigate the map $g_{2k-1}\colon C_{2k-1}^{M}\to C_{2k-1}^{M}$. It is represented by a $2\times 2$ matrix with entries in $\mathbb{Z}G$, with respect to the basis corresponding to $E_{2k-1}$ and $E_{2k-1}^{\prime}$. Since $g_{2k-1}\big{|}_{\mathbb{Z}G\oplus\left\{0\right\}}=\operatorname{Id}$, the matrix is $\left(\begin{smallmatrix}1&0\\ x&y\end{smallmatrix}\right)$ for some $x,y\in\mathbb{Z}G$. To compute $x$ and $y$, we consider the difference between $g$ and $\operatorname{Id}$ on $E_{2k-1}^{\prime}$. Since $g\big{|}_{\partial E^{\prime}_{2k-1}}=\operatorname{Id}$, the cell $E^{\prime}_{2k-1}$ and $g(E^{\prime}_{2k-1})$ together form a map $S^{2k-1}\rightarrow S^{1}\times L$. This map represents an element of $\pi_{2k-1}(S^{1}\times L)\cong\pi_{2k-1}(\mathbb{R}\times S^{2k-1})\cong\mathbb{Z}$, so its homotopy class is an integer multiple of that of the inclusion $S^{2k-1}\rightarrow\mathbb{R}\times S^{2k-1}$, composed with the projection ×$\mathbb{R}\times S^{2k-1}\rightarrow S^{1}\times L$. The image of the fundamental class of $S^{2k-1}$ under this composition is represented by the chain $\Sigma\in\mathbb{Z}G\oplus\left\{0\right\}<\mathbb{Z}G^{2}\cong C^{M}_{2k-1}$. So $E^{\prime}_{2k-1}$ and $g(E^{\prime}_{2k-1})$ differ by some integer multiple of $\Sigma\in\mathbb{Z}G\oplus\left\{0\right\}$. Therefore (by changing $g$ on $E^{\prime}_{2k-1}$ and $E^{\prime}_{2k}$, to replace it with another cellular map which is homotopic to it) we can assume that $y=1$ and $x=a\Sigma$ for some $a\in\mathbb{Z}$. That is, $g_{2k-1}=\left(\begin{smallmatrix}1&0\\ a\Sigma&1\end{smallmatrix}\right)$.
Since $(g_{j})$ is a chain map, we have $d^{M}_{2k}\circ g_{2k}=g_{2k-1}\circ d^{M}_{2k}$. When converted to matrices this yields
$g_{2k}(1-\alpha,\beta^{r_{k}}-1)=(1-\alpha,\beta^{r_{k}}-1)\left(\begin{smallmatrix}1&0\\ a\Sigma&1\end{smallmatrix}\right)=(1-\alpha,\beta^{r_{k}}-1),$
since $\beta^{r_{k}}\Sigma=\Sigma$. Looking at the first coordinate, we have $g_{2k}(1-\alpha)=(1-\alpha)$. Multiplication by $(1-\alpha)$ is an injective map $\mathbb{Z}G\rightarrow\mathbb{Z}G$, so we deduce that $g_{2k}=1$.
Therefore $g_{j}$ is an isomorphism for every $j$. By Lemma 2.18 we can compute the torsion via the formula $\tau(g)=\sum_{j=0}^{2k}(-1)^{j}[g_{j}]$. Since $g_{2k-1}$ is an elementary matrix and $g_{j}=\operatorname{Id}$ for $j\neq 2k-1$, this implies that $\tau(g)=0$. ∎
Lemma 6.9.
For every homotopy automorphism $f\colon S^{1}\times L\rightarrow S^{1}\times L$ there is another homotopy automorphism $g$ such that $\tau(f)=\tau(g)$ and $g\big{|}_{L\cup\operatorname{sk}_{2k-2}(L\times S^{1})}=\operatorname{Id}$.

### v3

@@ 5.2. Bass-Heller-Swan decomposition of $\operatorname{Wh}(C_{\infty}\times G)$
Let $G$ be a group and let $n\in\mathbb{Z}$. The decomposition of Theorem 5.9 restricts to the following isomorphisms:
@@
5.3. The groups $\mathcal{J}_{n}(C_{\infty}\times C_{m})$, $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$ for $n$ even
@@
By Proposition 5.10 and Theorem 5.7, for any group $G$ and any $n$, if $NK_{1}(\mathbb{Z}G)\neq 0$, then $\mathcal{J}_{n}(C_{\infty}\times G)$ and $\mathcal{I}_{n}(C_{\infty}\times G)$ are not finitely generated. If $G$ is finite and $n$ is even, then we also have the following.
@@
The next two theorems are the main results of Section 5. They are established as a consequence of Theorems 11.1 and 11.2 respectively (see Section 11).
@@
Suppose that $n$ is even and $m\geq 2$ is an integer. Then
@@
    $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=\infty$ if and only if $|\mathcal{I}_{n}(C_{\infty}\times C_{m})|=\infty$ if and only if $m$ is not square-free;
  4. (iv)
    $|\mathcal{I}_{n}(C_{\infty}\times C_{m})|\to\infty$ super-exponentially in $m$, and hence we also have that
@@
Part (i) says that the $m\geq 2$ for which $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=1$ are as follows:
[…]
Part (ii) says that $|\mathcal{I}_{n}(C_{\infty}\times C_{m})|=1$ if and only if $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=1$ or $m\in\{15,29\}$.
Proof.
First note that (iii) immediately follows from Theorem 5.8 and Lemma 5.12. For the proofs of (i), (ii) and (iv), we can therefore restrict to the case where $m$ is square-free, when $NK_{1}(\mathbb{Z}C_{m})=0$. By Propositions 5.10 and 5.13 we then have:
@@
The results now follow from Theorem 11.1. ∎
@@
$\liminf_{m\to\infty}|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|=1,\quad\limsup_{m\to\infty}|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))|=\infty.$
@@
Hence parts (i) and (iii) follows from Theorem 11.2, which are proved in Section 11. Part (ii) follows from the fact that $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ is finite [Sw60] (see also Proposition 8.4 (ii)). ∎
@@ 6. Homotopy automorphisms of $S^{1}\times L$
Fix integers $k,m\geq 2$ and $q_{1},\ldots,q_{k}$ such that $\gcd(m,q_{j})=1$. Let $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ be a (2k1)(2k-1)-dimensional lens space with $\pi_{1}(L)\cong C_{m}$. We are interested in the $2k$-dimensional product manifold $S^{1}\times L$ and its homotopy automorphisms, which are given as follows.
Lemma 6.1 ([Kh17, Corollary 6.2]).
We have
$\operatorname{hAut}(S^{1}\times L)\cong\pi_{1}(\operatorname{Map}(L))\rtimes(\operatorname{hAut}(S^{1})\times\operatorname{hAut}(L))$
where $\operatorname{Map}(L)$ denotes the space of continuous maps $L\rightarrow L$ (with basepoint $\operatorname{Id}_{L}$). In particular, $\operatorname{hAut}(S^{1}\times L)$ is generated by maps of the form $f\times g$ and $r_{h}$, where $f\in\operatorname{hAut}(S^{1})$, $g\in\operatorname{hAut}(L)$, $h\colon S^{1}\rightarrow\operatorname{Map}(L)$ and $r_{h}(x,y)=(x,h(x)(y))$.
We can deduce that every homotopy automorphism of $S^{1}\times L$ is simple.
@@
Proof.
Since $S^{1}$ and $L$ are odd dimensional, it follows from Corollary 2.33 that homotopy equivalences of the form $f\times g$ have vanishing Whitehead torsion. If $h\colon S^{1}\rightarrow\operatorname{Map}(L)$ is any pointed map, then $r_{h}$ is a fibre homotopy equivalence covering the identity of $S^{1}$, and acting as the identity on the copy of $L$ that is the fibre over the basepoint:
[…]
Hence $\tau(r_{h})=0$ by Proposition 2.34, and because $\tau(\operatorname{Id})=0$. ∎
Let $G=C_{\infty}\times C_{m}\cong\pi_{1}(S^{1}\times L)$. Our next goal is to determine which automorphisms of $G$ can be realised by homotopy automorphisms of $S^{1}\times L$, i.e. to describe the image of the map $\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(G)$. The automorphisms of $G=C_{\infty}\times C_{m}$ can be expressed as matrices of the form $\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)$, where $a\in\{\pm 1\}$, $b\in C_{m}$ and $c\in C_{m}^{\times}\cong\operatorname{Aut}(C_{m})$. In particular, we have $\operatorname{Aut}(G)\cong C_{m}\rtimes(C_{2}\times C_{m}^{\times})$.
Recall from Lemma 6.1 that the generators of $\operatorname{hAut}(S^{1}\times L)$ are of the form $f\times g$ or $r_{h}$. The induced automorphisms of $G$ are given by $\pi_{1}(f\times g)=\left(\begin{smallmatrix}\pi_{1}(f)&0\\ 0&\pi_{1}(g)\end{smallmatrix}\right)$ and $\pi_{1}(r_{h})=\left(\begin{smallmatrix}1&[\operatorname{ev}\circ h]\\ 0&1\end{smallmatrix}\right)$, where $\operatorname{ev}\colon\operatorname{Map}(L)\rightarrow L$ is given by evaluation at the basepoint $y_{0}\in L$, so that $(\operatorname{ev}\circ h)(x)=h(x)(y_{0})$.
Both elements of $\{\pm 1\}\cong\operatorname{Aut}(C_{\infty})$ can be realised by homotopy automorphisms (or even diffeomorphisms) of $S^{1}$. The subgroup of realisable automorphisms of $C_{m}$ depends on the dimension (2k1)(2k-1) of the lens space $L$ as follows.
Lemma 6.3 ([Co73, Statement 29.5]).
Let $c\in C_{m}^{\times}$. There is a homotopy automorphism $g\colon L\rightarrow L$ such that $\pi_{1}(g)=c$ if and only if $c^{k}\equiv\pm 1\,\,\text{mod}\,m$.
Finally, we have the following.
@@
There is a map $h\colon S^{1}\rightarrow\operatorname{Diff}(L)\subset\operatorname{Map}(L)$ such that $\operatorname{ev}\circ h]=1\in C_{m}\cong\pi_{1}(L)$. Hence the resulting diffeomorphism $r_{h}\in\operatorname{Diff}(S^{1}\times L)$ has $\pi_{1}(r_{h})=\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)$.
@@
Let $S^{2k-1}=\{(z_{1},\ldots,z_{k})\mid\sum|z_{j}|^{2}=1\}\subseteq\mathbb{C}^{k}$. Recall that $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ is the quotient of $S^{2k-1}$ by the $C_{m}$-action generated by $(z_{1},\ldots,z_{k})\mapsto(\zeta^{q_{1}}z_{1},\ldots,\zeta^{q_{k}}z_{k})$, where $\zeta=e^{2\pi i/m}$. Define $h\colon S^{1}\to\operatorname{Diff}(L)$ by $e^{2\pi it}\mapsto([z_{1},\ldots,z_{k}]\mapsto[\zeta^{q_{1}t}z_{1},\ldots,\zeta^{q_{k}t}z_{k}])$. Then the loop $\operatorname{ev}\circ h\colon S^{1}\rightarrow L$ is the standard generator of $\pi_{1}(L)\cong C_{m}$. ∎
Definition 6.5.
For a positive integer $k$ and $G=C_{\infty}\times C_{m}$, let $A_{2k}(m)\leq\operatorname{Aut}(G)$ denote the subgroup of matrices $\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)$ such that $c^{k}\equiv\pm 1$ mod $m$.
Theorem 6.6.
Let $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ be a (2k1)(2k-1)-dimensional lens space with $\pi_{1}(L)\cong C_{m}$. Then $\operatorname{Im}(\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(G))=A_{2k}(m)$.